3 times the complement of an angle is 30 more than the supplement of the original angle. What is the measure of the angle
Answer by sachi(548)

3 times the complement of an angle is 30 more than the supplement of the original angle.
so if the angle is x then its complement is 90-x & supplement is 180-x
so 3(90-x)=(180-x)+30
or 270-3x=210-x
or 3x-x=270-210
or 2x=60
or x=30 the original angle